# SAMG: State-Action-Aware Offline-to-Online Reinforcement Learning with Offline Model Guidance

### v1

@@ Introduction
By integrating the C-VAE form distribution-aware weights into Equation 1, the following practical updating expression can be obtained:
$Q(s,a)=$ $r(s,a)+\gamma\left[(1-p_{o}(s,a))Q(s^{\prime},a^{\prime})\right.$ (4)
$+\left.{\color[rgb]{0,0,1}{p_{o}(s,a)Q^{off}(s^{\prime},a^{\prime})}}\right].$
Comprehension of distribution-aware weight. The distribution-aware weight module actually attempts to depict the characteristics of the complex distribution represented by the offline dataset. Considering the high-dimensional and continuous property of the state-action data, it is challenging to directly extract the probability characteristics from the state-action pair. Therefore, it is required to perform dimension reduction of the distribution of the offline dataset, i.e., the complex distribution is simplified to be a normal distribution.
Analysis of SAMG
Intuitive analysis of SAMG
Intrinsic reward form of SAMG. Equation 4 can be derived as below:
$Q(s,a)=$ $\left[r(s,a)+\gamma p_{o}(s,a)(Q^{off}(s^{\prime},a^{\prime})-Q(s^{\prime},a^{\prime}))\right]$ (5)
$+\gamma Q(s^{\prime},a^{\prime}).$
Equation 5 indicates that the induced offline information could be treated in the form of intrinsic reward. The intrinsic reward is composed of the difference between offline and online state-action value factored by the distribution-aware weight. The intrinsic reward term can be analyzed in two possible situations. If the state-action pair is within the offline distribution where the offline Q-value is reasonably trained and the distribution-aware wight $\alpha$ is notable, then this term suggests that the higher the offline Q-value, the higher the reward. Hence, this intrinsic reward term encourages state-action pairs with higher performance. However, if the state-action pair falls outside the offline distribution where the offline Q-value may be mis-estimated, the wight $p_{o}$ is paltry or even set to zero according to Equation 3 and this term is insignificant. Therefore, SAMG could properly utilize the knowledge of the offline model, which is exactly what we expect.
Bellman form of SAMG. Furthermore, Equation 1 can be derived to the following equation by applying the bellman operator:
$Q(s,a)=(1-p_{o}(s,a))\mathcal{B}^{\pi}Q+p_{o}(s,a)\mathcal{B}^{\pi}Q^{off}.$ (6)
where $\mathcal{B}^{\pi}Q$ is the standard bellman equation and $\mathcal{B}^{\pi}Q^{off}=r(s,a)+\gamma Q^{off}(s^{\prime},a^{\prime})$ is the offline bellman equation of the fixed offline state action function.
It can be obtained that our method actually incorporates the standard Bellman equation with the offline Bellman equation by weight $p_{o}(s,a)$ and this paradigm just combines the offline opinion with the online opinion weighted by the sample property. More importantly, SAMG reverts to the vanilla algorithm when dealing with OOD samples. This allows SAMG to take full advantage of the conservative settings of the vanilla algorithm for OOD samples. Therefore, SAMG does not induce over-estimation on OOD samples even though SAMG gets rid of the offline dataset. Consequently, SAMG effectively utilizes the vanilla wisdom for handling OOD samples.
Theoretical analysis of SAMG
In this section, the temporal difference algorithm (Sutton 1988; Haarnoja et al. 2018) is adopted and Equation 4 is proven to still converge to the same optimalily, even with an extra term induced in the tabular setting. For the theoretical tools, SAMG gets rid of the offline dataset and therefore diverges from the hybrid realm of Song et al. (Song et al. 2023) and offline RL scope limited by the dataset, but aligns completely with traditional online RL algorithms (T. Jaakkola and Singh 1994; Thomas 2014; Haarnoja et al. 2018).
Contraction property. Since the standard Bellman operator is broken and a term that could be considered as offline Bellman operator is introduced, we first proof that SAMG still satisfies the contraction property (Keeler and Meir 1969). The contraction property forms the basis for the necessity of convergence of the iterative RL algorithms. The related theorem and detailed proof can be found in Appendix A.1.
Convergence optimality and estimation error. Formally, the estimated state-value function at time-step $k$ for any given pair $(s,a)$ is denote by $Q_{k}(s,a)$ and the iterative TD updating form of Equation 4 is shown as below:
$\begin{split}&Q_{k+1}(s,a)-Q_{k}(s,a)\\ =&\alpha_{k}(s,a)\left[Q_{k}(s,a)-\left(r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\right)\right]-\\ &\alpha_{k}(s,a)\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right)\end{split}$ (7)
To prove the optimality of SAMG, the Dvoretzky’s Theorem (Dvoretzky 1959) is introduced.
Lemma 1 (Dvoretzky’s Theorem).
(a) $\sum_{k}\alpha_{k}(s)=\infty,$ $\sum_{k}\alpha_{k}^{2}(s)<\infty$, $\sum\limits_{k}\beta_{k}^{2}(s)<\infty$, $\mathbb{E}[\beta_{k}(s)|\mathcal{H}_{k}]\leq\mathbb{E}[\alpha_{k}(s)|\mathcal{H}_{k}]$ uniformly almost surely;
(b) $\big{\|}{\mathbb{E}[\eta_{k}(s)|\mathcal{H}_{k}]}\big{\|}_{\infty}\leq\gamma\big{\|}\Delta_{k}\big{\|}_{\infty}$, with $\gamma\in(0,1)$;
(c) $var\left[\eta_{k}(s)|\mathcal{H}_{k}\right]\leq C\left(1+\big{\|}\Delta_{k}(s)\big{\|}_{\infty}\right)^{2}$, with $C$ a constant. Here, $\mathcal{H}_{k}=\{\Delta_{k},\Delta_{k-1},\cdots,\eta_{k-1},\cdots,\alpha_{k-1}$, $\cdots,\beta_{k-1},\cdots\}$ denotes the historical information. The term $\|\cdot\|_{\infty}$ represents the maximum norm.
The incremental form of SAMG is derived and proven to satisfy these conditions and the following optimality conclusion is obtained. Details refer to Appendix A.2.
Theorem 1 (Optimality of SAMG).
Given a policy $\pi$, by the TD updating paradigm, $Q_{k}(s,a)$ of SAMG converges almost surely to $Q_{\pi}(s,a)$ as $k\to\infty$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ if $\sum_{k}\alpha_{k}(s,a)=\infty$ and $\sum_{k}\alpha_{k}^{2}(s,a)<\infty$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$.
Moreover, the following theorem on convergence estimation error is obtained during the derivation of the second condition.
Theorem 2 (Estimation error of SAMG).
The error term $\eta_{k}(s,a)$ is limited by,
$\big{\|}\mathbb{E}[\eta_{k}(s,a)]\big{\|}_{\infty}$
@@
where $10<\gamma_{\mathcal{F}}<$ 1 denotes the convergence coefficient of offline algorithm class $\mathcal{F}$ and $\mathcal{O}(\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty})$ denotes the high-order infinitesimal of $\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}$.
@@
Theorem 2 indicates that SAMG converges faster than standard RL algorithms on samples within the distribution of offline datasets, illustrating the distribution-aware advantage of SAMG, and at least the same speed on OOD samples. For samples within the distribution, the convergence speed depends on the offline confidence implied by $p(s,a)$, i.e., a higher $p(s,a)$ indicates a higher degree of offline-ness, corresponding to a smaller error term constrained by the term $(1-p(s,a))$ and ensuring faster convergence. For the OOD samples, the algorithm degenerates into the traditional algorithm because $p(s,a)$ is set to zero as stated in Equation 3. The theoretical result is highly consistent with the analysis of the expected performance as shown in Section Intuitive analysis of SAMG and demonstrates evident performance enhancement. Furthermore, the theoretical result indicates that the extent of algorithm improvement is determined by the sample coverage rate of the offline dataset. Specifically, the algorithm’s performance is enhanced with more comprehensive sample coverage, whereas its improvement is constrained in scenarios with limited sample diversity.
Experimental Results
Our experimental evaluation focuses on the performance of SAMG after fine-tuning based on four state-of-the-art algorithms on a wide range of offline benchmark tasks containing D4RL (Fu et al. 2020). All experiments are conducted with compute source detailed in Appendix D.
Baselines. (i) SAMG algorithms, the SAMG paradigm is constructed on a variety of state-of-the-art O2O RL algorithms, including CQL (Kumar et al. 2020), AWAC (Nair et al. 2020), IQL (Kostrikov, Nair, and Levine 2022)and Cal_QL (Nakamoto et al. 2023). All algorithms are implemented based on library CORL (Tarasov et al. 2024) with implementation details in Appendix C.1 and C.2. (ii) O2O RL algorithms, we implement the aforementioned O2O RL algorithms (CQL, AWAC, IQL and Cal_QL). We also implement TD3+BC (Fujimoto and Gu 2021) and SPOT (Wu et al. 2022). (iii) Behavior Cloning (BC), we implement Behavior cloning algorithms based on (Chen et al. 2020).
Benchmark tasks. We evaluate our algorithm and the baselines across multiple benchmark tasks: (1) The D4RL locomotion tasks (Fu et al. 2020), including three different kinds of environments (Halfcheetah, Hopper, Walker2d) where different robots are manipulated to complete different tasks on three different levels of datasets. (2) The AntMaze tasks that an "Ant" robot is controlled to explore and navigate to random goal locations in six levels of environments, as detailed in Appendix C.3.
@@
Table 1: Performance comparison. The D4RL normalized score (Fu et al. 2020) is evaluated of standard base algorithms (including CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2022), AWAC (Nair et al. 2020) and Cal_QL (Nakamoto et al. 2023), denoted as "Vanilla") in comparison to the base algorithms augmented with SAMG (referred to as "Ours"), as well as three baselines (BC, TD3BC (Chen et al. 2020), SPOT (Wu et al. 2022)). The superior scores are highlighted in blue.
Dataset1 CQL AWAC IQL Cal-QL BC TD3BC SPOT
Vanilla Ours Vanilla Ours Vanilla Ours Vanilla Ours
Hopper-mr 100.9(0.1) 103.7(1.3) 101.7(2.5) 110.0(1.2) 99.6(1.3) 101.8(0.9) 103.4(0.4) 108.7(0.1) 59.9 (2.2) 99.4(1.3) 99.2(4.6)
Hopper-m 72.9(6.9) 93.6(4.8) 102.3(2.4) 104.1(10.0) 70.9(3.7) 74.1(5.3) 88.6(7.6) 112.9(2.5) 58.1(1.7) 66.4(1.6) 87.1(5.3)
Hopper-me 111.7(0.5) 113.0(0.3) 112.0(1.8) 112.8(7.2) 85.1(28.7) 95.7(6.1) 102.1(0.2) 112.3(1.5) 83.4(14.6) 97.4(9.1) 97.8(6.5)
Antmaze-u 98.0(2.0) 99.0(1.0) 70.0(40.4) 87.0(13.2) 94.0(1.0) 95.6(2.0) 99.8(0.4) 97.0(1.1) 54.2(3.7) 72.4(28.5) 99.5(0.5)
Antmaze-ud 95.0(1.7) 93.0(3.2) 16.0(35.3) 65.0(7.1) 75.6(20.4) 94.0(4.6) 95.4(3.0) 96.4(1.8) 46.3(3.8) 46.2(14.5) 92.5(3.7)
Antmaze-md 88.0(7.9) 97.0(1.0) 0.0(0.0) 0.0(0.0) 90.6(3.8) 96.5(1.9) 97.0(2.8) 96.0(2.1) 0.0(0.0) 0.2(0.2) 92.0(1.5)
Antmaze-mp 90.0(4.6) 96.2(4.0) 0.0(0.0) 0.0(0.0) 92.0(3.8) 94.8(1.6) 98.0(1.2) 96.0(1.6) 0.8(0.6) 0.3(0.3) 97.2(1.7)
Antmaze-ld 71.6(1.4) 70.8(2.8) 0.0(0.0) 0.0(0.0) 67.4(4.2) 81.5(7.9) 85.6(9.4) 87.2(4.7) 0.0(0.0) 0.0(0.0) 85.5(3.2)
Antmaze-lp 61.2(0.7) 75.0(2.1) 0.0(0.0) 0.0(0.0) 68.6(6.2) 74.8(8.4) 91.0(4.6) 90.0(1.2) 0.0(0.0) 0.0(0.0) 81.0(16.8)
@@
    mr: medium-replay, m: medium, l: large, me: medium-expert, d: diverse, p: play.
Figure 2: Algorithm performance in the online fine-tuning phase after offline pretraining. This figure plots the performance of SAMG and the other baselines during the online fine-tuning phase on part of the benchmarks (halfcheetah and walker2d). SAMG outperforms or at least matches the previous methods. The results are averaged on 5 seeds.
Empirical results
We first represent the normalized scores of the vanilla algorithms with and without SAMG integrated. The normalized scores through training on Walker2d and Halfcheetah environments are shown in Figure 2. Constrained by the space, we present the final normalized scores of the other environments (Hopper and Antmaze) in Table 1.
As shown in the Figure 2 and the Table 1, SAMG consistently outperforms the vanilla algorithms in the majority of environments, illustrating the superiority of our paradigm. Moreover, it is interesting to note that SAMG exhibits the best performance on the simple algorithm AWAC. SAMG still yields significant improvements on the other algorithms with more complex settings, but not as significant as AWAC. The reason for this counter-intuitive phenomenon is discussed in Appendix C.4., which just illustrates the effectiveness of SAMG. We present the cumulative regret on environment Antmaze in appendix C.5., further demonstrating the outstanding online sample efficicy of SAMG.
Although SAMG performs well in most environments, it is still worthwhile to notice SAMG may occasionally behave unsatisfactory(e.g., IQL-SAMG on Halfcheetah-medium-expert environment). We discuss in the Appendix C.6. that this exception is caused by the special property of the environment and IQL, rather than the defect of SAMG. We notice that the AWAC algorithm performs poorly in the Antmaze environment which also results in SAMG struggling to initiate. This is because AWAC algorithm is relatively simple and not competent for the complex environment of Antmaze, it is an inherent limitation of the AWAC algorithm, rather than an issue with SAMG. In summary, SAMG demonstrates performance increase by average 10.01% across four vanilla algorithms (for a total examined 15 tasks), indicating its robustness and versatility.
Figure 3: Ablation study for the distribution-aware weight.
Figure 4: Sensitivity test for the threshold $p_{om}$.

### v2

@@ 3.1 Offline-model-guidance Paradigm
$+\left.{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{p(s,a)Q^{off}(s^{\prime},a^{\prime})}}\right].$
where $Q(s,a)$ represents the estimated Q-function and $p\in(0,1)$ denotes a function class that gives a state-action-conditional coefficient and could be implemented with any reasonable form. The novel parts of the equation compared to the standard Bellman equation are marked in blue.
3.2 State-action-conditional Coefficient
State-action-conditional coefficient is proposed to address Problem 2. Intuitively, we tend to allocate higher values to samples within the offline distribution, as these samples are well-characterized and thoroughly pre-trained. Conversely, we have limited knowledge about samples distant from the offline distribution (treated as OOD samples), so lower values are appropriate. In summary, the state-action-conditional coefficient should capture the offline confidence of online samples. Any structure that satisfies the criteria can serve as an instantiation of $p(s,a)$.
The C-VAE model is primarily adopted to instantiate the state-action-conditional coefficient in this paper. C-VAE has been widely applied in estimating the behavior policies in offline RL with its ability to model complex distributions and provide sample probability.
Nevertheless, we observe that the previous C-VAE structure suffers from posterior collapse (Lucas et al., 2019; Wang et al., 2021), as illustrated in Appendix B.1.1. Posterior collapse implies that the encoder structure completely fails. The KL-divergence loss vanishes to zero for any input, and the latent output is just a standard normal distribution (0 for the mean and 1 for the variance). Thus, the decoder structure takes noise $z\sim\mathcal{N}(0,1)$ as input and reconstructs samples all by itself. This phenomenon is highly detrimental in our setting because we need the latent output to calculate the state-action-conditional coefficient but the structure now fails to operate.
To mitigate the adverse impacts of posterior collapse, we extend the variational conditional information to include state-action pairs and reconstruct the next state from the decoder. This approach complicates the modeling process and develops a state-action-conditional structure. Additionally, we employ the KL-annealing technique (Bowman et al., 2015), as detailed in Appendix B.1.2. Formally, SAMG adopts the following evidence lower bound (ELBO) on the offline dataset to get the state-action-conditional C-VAE:
$\max\limits_{\psi_{1},\psi_{2}}\mathbb{E}_{z\sim\text{Enc}_{\psi_{1}}}$ $\bigl{[}\log\text{Dec}_{\psi_{2}}(s^{\prime}|z,s,a)\bigr{]}$ (2)
$-\beta D_{KL}\bigr{[}{\text{Enc}_{\psi_{1}}(z|s,a)}||{\text{Dec}_{\psi_{2}}(z)}\bigr{]}$
where $\text{Enc}_{\psi_{1}}(z|s,a)$ and $\text{Dec}_{\psi_{2}}(s^{\prime}|z,s,a)$ represent the encoder and decoder structure respectively; $\text{Dec}_{\psi_{2}}(z)$ denotes the prior distribution of the encoder; and $D_{KL}\left[p||q\right]$ denotes the KL divergence. The former error term denotes the reconstruction loss while the latter denotes the KL divergence between the encoder distribution and the prior distribution.
3.3 Coefficient Generation and Adaptive Updates
Posterior Distribution Generation
To evaluate the effectiveness of improved C-VAE structure, we input the offline dataset to the trained C-VAE model and record the mean and variance values of encoder output, as illustrated in the Appendix B.1.3. The results indicate that posterior collapse is significantly alleviated. However, it is still unreliable to directly utilize the latent information $z$ because it is sampled from the narrow normal distribution and the sampling randomness overshadows the distribution information. To address this issue, we use the deterministic outputs $(z_{m},z_{v})$ in place of less reliable $z$. Since the statistical distribution of the encoder outputs $(z_{m},z_{v})$ still resembles a normal distribution, we fit these outputs to the corresponding normal distribution to represent the offline dataset. Specifically, the mean $z_{m}$ is modeled as $\mathcal{N}_{M}(\mu_{m},\sigma_{m})$ and the standard $z_{v}$ is modeled as $\mathcal{N}_{S}(\mu_{v},\sigma_{v})$.
Then, the offline coefficient of a given sample $(x_{m},x_{v})$ can be devised as the probability falling within a distance less than $|z_{m}-\mu_{m}|$ for mean and $|z_{v}-\mu_{v}|$ for variance. The cumulative distribution function $F_{X}(x)$ is leveraged to get the probability of an online sample aligning with the offline distribution. The intermediate probability can be obtained:
@@ 3.3 Coefficient Generation and Adaptive Updates
where $\omega$ is the weight of mean and standard. This weight is set to 1 because the estimation error of the mean is significantly smaller than that of the standard.
Moreover, in cases where the sample diverges notably from the offline distribution, the information about the sample is unknown so the probability is manually set to be 0, with a hyperparameter $p^{vae}_{m}$ introduced. The eventual equation to calculate the probability $p^{off}$ of a given sample $(z_{m},z_{v})$ is illustrated below:
$\begin{split}p^{vae}(s,a)=\left\{\begin{aligned} &p^{int}(s,a),&p^{int}(s,a)\geq p^{vae}_{m}\\ &0,&p^{int}(s,a)<p^{vae}_{m}\end{aligned}\right.\end{split}$ (4)
By integrating the C-VAE form state-action-conditional coefficients into Eq. (1), the following practical updating equation can be obtained:
$Q(s,a)=$ $r(s,a)+\gamma\left[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-p^{off}(s,a))}Q(s^{\prime},a^{\prime})\right.$ (5)
$+\left.{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{p^{off}(s,a)Q^{off}(s^{\prime},a^{\prime})}}\right].$
Comprehension of state-action-conditional coefficient. The state-action-conditional coefficient module actually attempts to depict the characteristics of the complex distribution represented by the offline dataset. Considering the high-dimensional and continuous property of the state-action data, it is challenging to directly extract the probability characteristics from the state-action pair. Therefore, it is required to perform dimension reduction of the distribution of the offline dataset, i.e., the complex distribution is simplified to be a normal distribution.
Adaptive Coefficient Updates
During the online training process, the fine-tuned model gradually adapts to the environment and learns to handle some of the OOD samples. However, the offline guidance and state-action-conditional coefficient remain fixed, which becomes more and more unsuitable for probability generation. To address this issue (Problem 3), adaptive coefficient updates are proposed. At fixed intervals, OOD samples from the current period are collected, and these OOD samples with the lowest offline Q-loss are identified and reclassified as mastered OOD samples. They are then used to retrain the VAE model, allowing the VAE coefficient to adapt to the online fine-tuning process. Meanwhile, the offline pre-trained model is updated to the Q-function at these specific intervals. Refer to Appendix B.1.4 for complete implementation.
4 Analysis of SAMG
4.1 Intuitive analysis of SAMG
Intrinsic Reward Analysis highlights the importance of probability-based coefficient. Specifically, Eq. (5) can be derived as below:
$Q(s,a)=\bigl{[}r(s,a)+r^{in}(s,a)\bigr{]}+\gamma Q(s^{\prime},a^{\prime}).$ (6)
where $r^{in}(s,a)=\gamma p(s,a)(Q^{off}(s^{\prime},a^{\prime})-Q(s^{\prime},a^{\prime}))$. Eq. (6) indicates that the induced offline information could be treated as intrinsic reward.
Previous work has revealed that intrinsic reward may cause training instability or even algorithm degradation (Chen et al., 2022; McInroe et al., 2024). However, the intrinsic reward form of SAMG is reasonable and stable thanks to the probability-shape coefficient.
Specifically, the intrinsic reward term describes the difference between offline and online Q-values, weighted by the state-action-conditional coefficient. It can be analyzed in two scenarios. Firstly, if the state-action pair lies within the offline distribution, where the offline Q-value is well trained and the state-action-conditional coefficient $\alpha$ is significant. In this case, this term suggests that higher offline Q-values correspond to higher potential returns. Hence, it encourages exploring state-action pairs with higher performance. Conversely, if the state-action pair falls outside the offline distribution, where the offline Q-value may be erroneously estimated. This term becomes negligible or is even set to zero, as specified in Eq. (4). Therefore, it can filter out inaccurate and questionable information. In summary, SAMG is able to properly retain the offline knowledge without introducing inappropriate intrinsic rewards.
4.2 Theoretical analysis of SAMG
In this section, we adopt the temporal difference paradigm (Sutton, 1988; Haarnoja et al., 2018) in the tabular setting and prove that Eq. (5) still converges to the same optimality, even with an extra term induced. For the theoretical tools, SAMG gets rid of the offline dataset and therefore diverges from the hybrid realm of Song et al. (Song et al., 2023) and offline RL scope limited by the dataset, but aligns with traditional online RL algorithms (T. Jaakkola and Singh, 1994; Thomas, 2014; Haarnoja et al., 2018).
Contraction Property is considered and proven to still hold since the standard Bellman operator is broken (Keeler and Meir, 1969). The related theorem and detailed proof can be found in Appendix A.1.
Convergence Optimality. Formally, the iterative TD updating form of Eq. (5) is demonstrated as below:
$\begin{split}&Q_{k+1}(s,a)-Q_{k}(s,a)=\\ &\alpha_{k}(s,a)\left[Q_{k}(s,a)-\left(r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\right)\right]-\\ &\alpha_{k}(s,a)\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right)\end{split}$ (7)
where the estimated state-value function at time-step $k$ for given $(s,a)$ pair is denoted as $Q_{k}(s,a)$. The learning rate at time-step $k$ is represented as $\alpha_{k}$. For simplicity, the state-action conditional coefficient $p^{off}(s,a)$ and threshold $p^{off}_{m}(s,a)$ are denoted as $p(s,a)$ and $p_{m}(s,a)$, respectively.
The proof pipeline involves deriving the incremental form of SAMG  then proving that the incremental form satisfies Dvoretzky’s Theorem (Dvoretzky, 1959) and finally arriving at the following theorem. Details refer to Appendix A.2.
Theorem 4.1 (Convergence property of SAMG).
For a given policy $\pi$, by the TD updating paradigm, $Q_{k}(s,a)$ of SAMG converges almost surely to $Q^{\pi}(s,a)$ as $k\to\infty$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$ if $\sum_{k}\alpha_{k}(s,a)=\infty$ and $\sum_{k}\alpha_{k}^{2}(s,a)<\infty$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$.
Convergence Speed. Moreover, the specific expression for the contraction coefficient is proven as follows, illustrating the faster convergence speed of SAMG. See Appendix A.3 for further details.
Theorem 4.2 (Convergence speed of SAMG).
The Bellman operator of SAMG satisfies the contraction property $\big{\|}\mathcal{B}(x)-\mathcal{B}(y)\big{\|}\leq\gamma_{o}\big{\|}x,y\big{\|}$ for all $x,y\in\mathcal{Q}$. $\mathcal{Q}$ represents the Q function space. The contraction coefficient of SAMG $\gamma_{o}(s,a)$ is bounded above by the following expression:
[…]
@@ 4.2 Theoretical analysis of SAMG
where $C=\big{\|}\Delta_{off}(s,a)\big{\|}_{\infty}\big{/}\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}$ denotes the ratio of the offline and online suboptimality bounds, $\big{\|}\Delta_{off}(s,a)\big{\|}_{\infty}$ denotes the offline suboptimality bound $\left[V^{*}(s)-V^{\pi_{off}}(s)\right]$, $\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}$ denotes the suboptimality bound of the k-th iteration of online fine-tuning and $10<\gamma_{\mathcal{F}}<$ 1 denotes the convergence coefficient of offline algorithm class $\mathcal{F}$.
@@
The upper equation of Theorem 4.2 holds for in-distribution samples, which are well mastered by the offline model. Therefore, the offline suboptimality bound is substantially tighter compared to the online bound. This illustrates that the offline model guidance significantly accelerates the online fine-tuning process by providing more accurate estimations for in-distribution samples. Specifically, for these samples, the convergence speed depends on the offline confidence implied by $p(s,a)$, i.e., a higher $p(s,a)$ indicates a higher degree of offline-ness, corresponding to a smaller error term constrained by the term $(1-p(s,a))$ and ensuring faster convergence. For the OOD samples, the algorithm degenerates into the traditional algorithm because $p(s,a)$ is set to zero as defined in Eq. (4). This theoretical result is highly consistent with the analysis of the expected performance as stated in Section 4.1. Furthermore, it indicates that the extent of algorithm improvement is influenced by the sample coverage rate of the offline dataset. Specifically, the offline guidance is more reliable with more complex sample coverage, whereas the guidance is constrained with limited sample diversity.
5 Experimental Results
Our experimental evaluations focus on the performance of SAMG during the online fine-tuning process based on three state-of-the-art algorithms within D4RL (Fu et al., 2020), covering diverse environments and task complexities, as detailed below.
@@ 5 Experimental Results
Table 1: Performance comparison. The D4RL normalized score (Fu et al., 2020) is evaluated of standard base algorithms (including CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2022) and AWAC (Nair et al., 2020), denoted as “Vanilla") in comparison to the base algorithms augmented with SAMG (referred to as “Ours"), as well as three baselines (TD3BC (Chen et al., 2020), SPOT (Wu et al., 2022), Cal_QL (Nakamoto et al., 2023) and EDIS (Liu et al., 2024)). The superior scores are highlighted in blue. The result is the average normalized score of 5 random seeds $\pm$ (standard deviation).
Dataset1 CQL AWAC IQL TD3BC SPOT Cal_QL EDIS
Vanilla Ours Vanilla Ours Vanilla Ours
Hopper-mr 100.6(1.8) 103.7(1.3) 99.4(1.3) 108.3(0.2) 86.2(16.1) 100.4(0.9) 64.4(21.5) 68.0(11.2) 80.9(38.2) 83.0(26.8)
Hopper-m 60.2(2.7) 88.3(6.0) 88.2(14.6) 102.5(1.8) 62.1(7.4) 68.4(2.9) 66.4(3.5) 54.6(7.1) 78.1(8.7) 30.1(8.9)
Hopper-me 110.8(1.0) 113.0(0.3) 101.9(20.5) 112.8(7.2) 103.5(8.7) 108.1(3.1) 101.2(9.1) 82.6(11.5) 109.1(0.2) 78.4(3.5)
Half-mr 48.0(0.5) 57.8(1.7) 48.9(1.1) 62.8(3.3) 45.1(0.6) 49.6(1.0) 44.8(0.6) 42.4 (3.7) 51.6(0.8) 82.9(1.2)
Half-m 47.6(0.2) 59.0(0.7) 54.2(1.1) 69.5(0.9) 49.3(0.1) 62.5(1.5) 48.1(0.2) 45.9(2.4) 63.2(2.5) 66.4(11.7)
Half-me 95.2(1.0) 97.2(0.8) 94.8(1.3) 96.7(1.0) 91.6(0.9) 82.3(11.7) 90.8(6.0) 87.4(7.4) 95.6(4.3) 90.2(1.4)
Walker2d-mr 82.7(0.7) 88.4(5.0) 93.8(3.4) 120.1(3.1) 87.1(3.3) 99.5(2.4) 85.6(4.0) 69.2 (6.2) 97.1(2.5) 46.9(23.6)
Walker2d-m 60.2(2.7 82.9(1.8) 87.8(0.8) 103.6(4.5) 83.4(1.6) 88.6(4.7) 82.7(4.8) 79.5(2.4) 83.6(0.8) 76.2(16.7)
Walker2d-me 109.5(0.5) 112.5(0.7) 112.7(0.9) 129.2(3.6) 113.6(1.1) 116.3(3.7) 110.0(0.4) 87.8(3.9) 110.7(0.4) 107.9(10.3)
Antmaze-u 92.0(1.7) 97.0(1.4) 70.0(40.4) 87.0(13.2) 83.3(6.1) 94.0(1.2) 70.8(39.2) 30.8(12.9) 96.8(0.4) 95.0(7.0)
Antmaze-ud 58.0(32.0) 62.3(12.4) 15.0(35.3) 75.0(7.0) 33.3(4.4) 77.8(0.8) 44.8(11.6) 44.8(6.5) 63.8(43.4) 72.5(32.5)
Antmaze-md 82.4(2.2) 89.1(3.5) 0.0(0.0) 0.0(0.0) 76.5(5.4) 96.5(1.9) 0.3(0.4) 36.3(11.0) 93.4(3.6) 82.4(4.8)
Antmaze-mp 85.6(6.6) 86.1(1.1) 0.0(0.0) 0.0(0.0) 76.2(4.6) 95.2(1.6) 0.3(0.4) 38.5(8.7) 94.0(2.2) 60.0(51.9)
Antmaze-ld 62.8(7.4) 63.9(5.9) 0.0(0.0) 0.0(0.0) 45.4(7.7) 81.5(7.9) 0.0(0.0) 0.0(0.0) 78.8(5.8) 32.5(15.0)
Antmaze-lp 55.0(8.4) 60.7(1.5) 0.0(0.0) 0.0(0.0) 48.8(7.7) 74.8(8.4) 0.0(0.0) 0.0(0.0) 73.0(19.4) 35.0(17.3)
Pen-b 90.0(4.6) 96.2(4.0) 63.3(39.7) 70.1(26.6) 86.7(24.6) 106.0(27.8) 6.4(4.37) 2.80(11.82) -0.03(4.10) 8.99(17.18)
Door-b -0.34(0.01) 70.8(2.8) 0.00(0.01) 7.29(3.14) -0.06(0.03) 13.08(3.20) -0.32(0.01) -0.16(0.05) -0.33(0.01) 0.09(0.03)
Relocate-b -0.28(0.12) 75.0(2.1) -8.84(1.22) -7.82(0.93) -0.01(0.04) 0.24(0.03) -0.21(0.01) -0.14(0.10) -0.31(0.03) -0.34(0.02)
@@
    Half: HalfCheetah, mr: medium-replay, me: medium-expert, d: diverse, p: play, u: umaze, ud: umaze-diverse, md: medium-diverse, mp: medium-play, ld: large-diverse, lp: large-play, b: binary.
Figure 3: Ablation analysis of the state-action-conditional coefficient.
Figure 4: Sensitivity test for the coefficient threshold $p^{vae}_{m}$.